What is the percent yield of copper when 13.8g of Cu2O reacts with Cu2S and 14.7 grams of copper is recovered?

2Cu2O + Cu2S → 6Cu + SO2
1
Expert's answer
2019-03-12T08:46:09-0400

2Cu2O + Cu2S"\to" 6Cu + SO2

m(Cu)="\dfrac{m(Cu2O)*6*M(Cu)}{2*M(Cu2O}=\dfrac{13.8g*6*63.546g}{2*143.09g/mol}=18.385g"

%yield="\dfrac{14.7g}{18.385g}*100=79.95"